Daniel is four years older than Miguel. In five years the sum of their age will be 30. What are their ages now?
Answer by macston(5194)

.
M=Miguel's age; D=Daniels age=M+4yrs
.
(M+5yrs)+(D+5yrs)=30
(M+5yrs)+(M+4yrs)+5yrs=30
2M+14yrs=30
2M=16yrs
M=8yrs ANSWER 1: Miguel is 8 years old.
D=M+4yrs=8yrs+4yrs=12yrs ANSWER 2: Daniel is 12 years old.
.